### v1

# When Does Multi-Task Learning Fail? Quantifying Data Imbalance and Task Independence in Metal Alloy Property Prediction

Abstract
Multi-task learning (MTL) has emerged as a promising approach in materials informatics, with the assumption that related material properties share underlying physical principles that can be leveraged to improve prediction accuracy. In this study, we critically examine this assumption by applying MTL to simultaneously predict three properties of metal alloys: electrical resistivity, Vickers hardness, and amorphous-forming ability. Using a union dataset of 54,028 samples from the AI-Hub Korea platform, we compare three approaches: independent single-task models, standard MTL with hard parameter sharing, and structured MTL with learned task relation graphs. Contrary to expectations, our results reveal a striking dichotomy: MTL significantly degrades regression performance (resistivity $R^{2}$: 0.897 $\rightarrow$ 0.844, $p<0.01$; hardness $R^{2}$: 0.832 $\rightarrow$ 0.694, $p<0.01$) while significantly improving classification performance (amorphous F1: 0.703 $\rightarrow$ 0.744, $p<0.05$; recall improved by 17%). Analysis of the learned task relation graphs reveals near-zero inter-task weights ($\sim$0.006), indicating that these alloy properties are fundamentally independent rather than coupled. We attribute the regression degradation to negative transfer caused by severe data imbalance (52,388 resistivity samples vs. 800 hardness samples), where the dominant task overwhelms minority tasks during joint optimization. These findings provide empirical guidelines for practitioners: use independent models for precise property regression, but leverage MTL for classification tasks where improved recall is critical for candidate screening. Our work highlights the importance of validating MTL assumptions before deployment in materials discovery pipelines.
keywords:
multi-task learning , negative transfer , metal alloys , property prediction , machine learning , materials informatics
journal: Computational Materials Science
[…]
[1]organization=Department of Electrical and Computer Engineering, Korea University, city=Seoul, postalcode=02841, country=Republic of Korea
1 Introduction
The accelerated discovery of advanced materials with tailored properties is a central goal of modern materials science. Machine learning (ML) has emerged as a powerful tool in this endeavor, enabling rapid prediction of material properties from composition and structure descriptors [schmidt2019recent, butler2018machine]. Among ML approaches, multi-task learning (MTL) has garnered significant attention for its promise to improve prediction accuracy by jointly learning multiple related properties [caruana1997multitask].
The appeal of MTL in materials science rests on a fundamental assumption: that different material properties share underlying physical mechanisms, and thus learning to predict one property can provide useful inductive bias for predicting others [caruana1997multitask]. For metal alloys, this assumption appears intuitive—electrical resistivity, mechanical hardness, and glass-forming ability are all influenced by atomic-level factors such as electronegativity differences, atomic size mismatch, and mixing enthalpy [zhang2014microstructures]. This has motivated numerous studies applying transfer learning [xie2018crystal, chen2021atomsets] and multi-task learning [debnath2025overcoming] to materials property prediction.
However, the question of when MTL actually helps—and when it may harm performance—has received insufficient attention in the materials informatics literature. The machine learning community has documented the phenomenon of “negative transfer,” where jointly learning unrelated or conflicting tasks degrades performance compared to independent learning [wang2019characterizing, standley2020tasks]. This is particularly concerning when datasets are severely imbalanced across tasks, as the optimization may be dominated by the largest task at the expense of smaller ones [chen2018gradnorm].
In this work, we conduct a systematic investigation of MTL for multi-property prediction in metal alloys, with a focus on understanding when and why MTL succeeds or fails. We make the following contributions:
  1. 1.
    We construct a comprehensive benchmark using the AI-Hub Korea metal alloy dataset, containing 54,028 unique samples with three target properties: electrical resistivity (52,388 samples), Vickers hardness (800 samples), and amorphous-forming ability (840 samples).
  2. 2.
    We implement and compare three modeling approaches: independent single-task neural networks, standard MTL with hard parameter sharing, and structured MTL with learned task relation graphs.
  3. 3.
    We provide unexpected but important findings: MTL significantly degrades regression performance while significantly improving classification performance. This challenges the common assumption that MTL universally benefits materials property prediction.
  4. 4.
    We analyze the root causes of this dichotomy through learned task relations (which reveal task independence) and gradient analysis (which reveals task interference from data imbalance).
  5. 5.
    We provide practical guidelines for materials scientists on when to apply MTL versus independent models.
The remainder of this paper is organized as follows. Section 2 reviews related work on MTL in materials science and the phenomenon of negative transfer. Section 3 describes our dataset, model architectures, and experimental setup. Section 4 presents comprehensive results with statistical analysis. Section 5 analyzes why MTL fails for regression but succeeds for classification, and provides practical recommendations. Section 6 summarizes our findings and outlines future directions.
2 Related Work
2.1 Multi-Task Learning in Materials Science
Multi-task learning has been increasingly applied to materials property prediction, motivated by the hypothesis that related properties share underlying physical principles. Xie and Grossman [xie2018crystal] demonstrated that graph neural networks can learn transferable representations for inorganic crystals, showing that pretraining on formation energy improves band gap prediction. Chen et al. [chen2021atomsets] developed AtomSets, a hierarchical transfer learning framework that leverages pre-trained representations on large datasets to improve predictions on smaller target datasets.
For molecular property prediction, Ramsundar et al. [ramsundar2015massively] showed that MTL on hundreds of biological assays can improve prediction of drug-target interactions. More recently, Wen et al. [wen2021bondnet] extended this to reaction property prediction, demonstrating benefits when tasks share chemical mechanisms.
However, most of these studies focus on scenarios where tasks are inherently related (e.g., multiple biological assays for the same target, or properties governed by similar physics). The question of whether arbitrary material properties benefit from joint learning has been less thoroughly examined.
2.2 Negative Transfer and Task Interference
The machine learning community has extensively studied the phenomenon of negative transfer, where knowledge transfer between tasks degrades rather than improves performance [wang2019characterizing]. Standley et al. [standley2020tasks] systematically analyzed task relationships in computer vision and found that many seemingly related tasks exhibit negative transfer when trained jointly.
Several factors contribute to negative transfer:
  • 1.
    Task dissimilarity: When optimal representations for different tasks conflict, shared parameters are forced into suboptimal compromises [crawshaw2020multi].
  • 2.
    Data imbalance: When tasks have vastly different sample sizes, gradient updates are dominated by the larger task, starving smaller tasks of learning signal [chen2018gradnorm].
  • 3.
    Gradient interference: Conflicting gradient directions across tasks can cause optimization instabilities [yu2020gradient].
Chen et al. [chen2018gradnorm] proposed GradNorm to dynamically balance gradient magnitudes across tasks, partially mitigating imbalance effects. Kendall et al. [kendall2018multi] introduced uncertainty-based task weighting. However, these methods address symptoms rather than the fundamental question of task compatibility.
2.3 Task Relation Learning
Rather than assuming task relationships a priori, recent work has explored learning task relations directly from data. Zamir et al. [zamir2018taskonomy] computed a “taskonomy” revealing which visual tasks transfer well to others. For molecular property prediction, several studies have learned task graphs indicating which properties provide useful auxiliary information for others [liu2022auto].
In materials science, explicit task relation learning remains underexplored. Most MTL studies either assume all properties are related (fully shared representations) or manually specify which properties should be grouped. Our work addresses this gap by learning task relations for metal alloy properties and using the learned structure to understand MTL performance.
2.4 MTL for Metal Alloys
Metal alloys present an interesting case study for MTL due to the complex interplay of their properties. High-entropy alloys (HEAs) in particular have attracted attention for their unusual property combinations [zhang2014microstructures]. Several studies have applied ML to predict HEA properties:
Debnath et al. [debnath2025overcoming] used MTL to predict yield strength and elongation of HEAs, finding modest improvements. Pei et al. [pei2020machine] predicted hardness and phase stability using composition features. However, these studies typically use small, curated datasets where all samples have all properties labeled—a scenario quite different from the realistic setting where different properties are measured on different sample sets.
Our work addresses this realistic scenario using a union dataset approach where each sample may have only a subset of properties labeled, employing masked loss functions to handle missing labels during training.
3 Methodology
3.1 Dataset
We utilize the metal alloy dataset from AI-Hub Korea, which contains experimental measurements of various properties for high-entropy alloys and related metallic systems. The dataset comprises three target properties:
  • 1.
    Electrical Resistivity ($\mu\Omega\cdotcm$): A continuous property measured for 52,388 samples.
  • 2.
    Vickers Hardness (GPa): A continuous property measured for 800 samples.
  • 3.
    Amorphous-Forming Ability: A binary classification (crystalline/amorphous) with 840 samples.
Critically, these properties were measured on largely non-overlapping sample sets. Traditional MTL approaches require all samples to have all labels, which would yield only a small intersection of samples. Instead, we employ a union dataset approach, concatenating all samples and using masked loss functions to handle missing labels. This yields a total of 54,028 samples for training.
3.1.1 Input Features
Each sample is represented by a 21-dimensional feature vector comprising:
  • 1.
    Elemental composition (15 features): Atomic fractions of Al, Ti, Cr, Fe, Co, Ni, Cu, Zr, Mo, W, Mn, Si, Mg, Re, and Ta.
  • 2.
[…]
All features were normalized to zero mean and unit variance based on training set statistics.
3.1.2 Data Splitting
We partition the data into training (70%), validation (15%), and test (15%) sets using stratified sampling to maintain property distributions. Table 1 summarizes the data statistics.
Table 1: Dataset statistics showing severe imbalance across properties.
Property Train Validation Test
Resistivity 36,670 7,859 7,859
Hardness 560 120 120
Amorphous 588 126 126
Total (Union) 37,818 8,105 8,105
The severe imbalance (65:1 ratio between resistivity and hardness) is a key factor in our analysis of MTL performance.
3.2 Model Architectures
We implement and compare three modeling approaches:
3.2.1 Independent Single-Task Models
For each property, we train an independent neural network with:
  • 1.
    Input layer: 21 features
  • 2.
    Hidden layers: 128 $\rightarrow$ 128 units with BatchNorm, ReLU, and Dropout (0.3)
  • 3.
    Output layer: 1 unit (linear for regression, sigmoid for classification)
3.2.2 Standard MTL with Hard Parameter Sharing
The standard MTL model shares a backbone across all tasks:
  • 1.
    Shared backbone: 21 $\rightarrow$ 128 $\rightarrow$ 128 with BatchNorm, ReLU, Dropout
  • 2.
    Task-specific heads: Each property has a 64-unit hidden layer followed by output
The loss function combines individual task losses with task-specific weights:
$\mathcal{L}_{MTL}=w_{r}\mathcal{L}_{res}+w_{h}\mathcal{L}_{hard}+w_{a}\mathcal{L}_{amor}$ (1)
where weights are set inversely proportional to dataset size ($w_{r}=1.0$, $w_{h}=65.0$, $w_{a}=62.0$) to mitigate imbalance.
3.2.3 Structured MTL with Task Relation Graphs
Building on standard MTL, we add a task relation graph module that learns the strength of information transfer between tasks. The module consists of:
  • 1.
    Task embeddings: Learnable 64-dimensional embeddings for each task
  • 2.
    Graph convolution: Two-layer GCN operating on a fully-connected task graph
  • 3.
    Edge predictor: MLP that predicts relation strength $w_{ij}\in[0,1]$ from task pairs
The learned task graph modulates auxiliary information flow:
$\hat{y}_{i}=f_{i}(z)+\alpha\sum_{j\neq i}w_{ij}\cdot g_{i}([\hat{y}_{j}])$ (2)
where $f_{i}$ is the primary prediction head, $g_{i}$ is an auxiliary fusion network, and $\alpha=0.1$ controls the auxiliary contribution.
The structured MTL loss includes regularization terms:
$\mathcal{L}_{Struct}=\mathcal{L}_{MTL}+\lambda_{1}\|W\|_{1}+\lambda_{2}(1-\text{tr}(W))$ (3)
encouraging sparsity in task relations ($\lambda_{1}=0.01$) and discouraging self-loops ($\lambda_{2}=0.1$).
3.3 Training Procedure
All models were trained using Adam optimizer with:
  • 1.
    Learning rate: 0.001 with ReduceLROnPlateau scheduling
  • 2.
    Batch size: 32
  • 3.
    Maximum epochs: 200
  • 4.
    Early stopping: Patience of 30 epochs based on validation loss
  • 5.
    Weight decay: $10^{-5}$
We handle missing labels using masked loss:
$\mathcal{L}_{task}=\frac{1}{\sum_{i}m_{i}}\sum_{i}m_{i}\cdot\ell(y_{i},\hat{y}_{i})$ (4)
where $m_{i}\in\{0,1\}$ indicates label availability and $\ell$ is MSE for regression or BCE for classification.
3.4 Evaluation Metrics
For regression tasks (resistivity, hardness):
  • 1.
    Root Mean Squared Error (RMSE)
  • 2.
    Mean Absolute Error (MAE)
  • 3.
    Coefficient of Determination ($R^{2}$)
For classification (amorphous):
  • 1.
    Accuracy
  • 2.
    F1-Score (harmonic mean of precision and recall)
  • 3.
    Area Under ROC Curve (AUC)
  • 4.
    Recall (sensitivity for detecting amorphous-forming alloys)
3.5 Statistical Testing
To ensure robustness, all experiments were repeated with 5 random seeds (42, 123, 456, 789, 1024). We report mean $\pm$ standard deviation and conduct paired t-tests to assess statistical significance of performance differences ($\alpha=0.05$).
4 Results
4.1 Main Performance Comparison
Table 2 presents the comprehensive comparison of all three modeling approaches across the three target properties.
Table 2: Performance comparison across modeling approaches. Best results for each metric are bolded. Statistical significance of Structured MTL vs. Independent is indicated: * $p<0.05$, ** $p<0.01$.
Property Metric Independent Standard MTL Structured MTL
Resistivity RMSE 0.252 $\pm$ 0.002 0.313 $\pm$ 0.019 0.310 $\pm$ 0.022**
MAE 0.161 $\pm$ 0.001 0.217 $\pm$ 0.018 0.215 $\pm$ 0.022**
$R^{2}$ 0.897 $\pm$ 0.001 0.842 $\pm$ 0.019 0.844 $\pm$ 0.021**
Hardness RMSE 0.407 $\pm$ 0.011 0.516 $\pm$ 0.075 0.548 $\pm$ 0.035**
MAE 0.312 $\pm$ 0.011 0.376 $\pm$ 0.069 0.396 $\pm$ 0.041**
$R^{2}$ 0.832 $\pm$ 0.009 0.724 $\pm$ 0.078 0.694 $\pm$ 0.038**
Amorphous Accuracy 0.810 $\pm$ 0.005 0.795 $\pm$ 0.025 0.819 $\pm$ 0.021
F1 0.703 $\pm$ 0.004 0.702 $\pm$ 0.050 0.744 $\pm$ 0.030*
AUC 0.923 $\pm$ 0.003 0.905 $\pm$ 0.018 0.912 $\pm$ 0.012
Recall 0.617 $\pm$ 0.011 0.670 $\pm$ 0.084 0.722 $\pm$ 0.052*
Key Finding 1: MTL significantly degrades regression performance. For resistivity prediction, independent models achieve $R^{2}=0.897$, while both MTL variants fall to $R^{2}\approx 0.84$—a 5.9% relative decrease. The degradation is even more pronounced for hardness, where $R^{2}$ drops from 0.832 (independent) to 0.694 (structured MTL)—a 16.6% relative decrease. All regression degradations are statistically significant at $p<0.01$.
Key Finding 2: MTL significantly improves classification performance. In stark contrast to regression, MTL improves amorphous classification. Structured MTL achieves the highest F1 score (0.744 vs. 0.703, $p=0.041$) and critically, the highest recall (0.722 vs. 0.617, p=0.013p=0.013)—a 17% relative improvement. This means MTL identifies 17% more amorphous-forming alloys.
Key Finding 3: Structured MTL provides no advantage over Standard MTL. Comparing Structured MTL to Standard MTL (Table 3), we find no statistically significant differences for any metric ($p>0.05$ for all comparisons). The task relation graph, despite its added complexity, does not improve over simple hard parameter sharing.
Table 3: Statistical comparison: Structured MTL vs. Standard MTL. No significant differences observed.
Property Metric $p$-value
Resistivity $R^{2}$ 0.865
Hardness $R^{2}$ 0.617
Amorphous F1 0.273
4.2 Learned Task Relations
Figure 1 shows the learned task relation graph from Structured MTL. Strikingly, all inter-task weights are near zero ($\sim$0.006 $\pm$ 0.012), indicating that the model learned no meaningful information transfer between tasks.
Figure 1: Learned task relation graph. All inter-task weights are near zero, indicating task independence.
Table 4 quantifies the learned relation strengths:
Table 4: Learned task relation strengths (mean $\pm$ std across seeds).
Relation Strength
Resistivity $\rightarrow$ Hardness 0.006 $\pm$ 0.012
Resistivity $\rightarrow$ Amorphous 0.006 $\pm$ 0.013
Hardness $\rightarrow$ Resistivity 0.007 $\pm$ 0.013
Hardness $\rightarrow$ Amorphous 0.006 $\pm$ 0.013
Amorphous $\rightarrow$ Resistivity 0.006 $\pm$ 0.013
Amorphous $\rightarrow$ Hardness 0.006 $\pm$ 0.012
This is a critical finding: despite the intuition that alloy properties should be related through shared physics, the data-driven approach reveals that these specific properties—as measured in this dataset—are fundamentally independent.
4.3 Training Dynamics Analysis
To understand why MTL degrades regression performance, we analyze training dynamics. Figure 2 shows validation loss curves for the three approaches.
Figure 2: Training curves showing validation loss over epochs. MTL models exhibit higher loss for regression tasks despite longer training.
For regression tasks, independent models converge to lower validation loss than MTL variants. This suggests that the shared representation learned by MTL is suboptimal for individual regression tasks—a hallmark of negative transfer.
4.4 Error Distribution Analysis
Figure 3 compares prediction error distributions across models. For hardness prediction, independent models show a tighter error distribution centered near zero, while MTL models exhibit heavier tails indicating more severe mispredictions.
Figure 3: Error distributions for each property. Independent models show tighter distributions for regression tasks.
4.5 Classification Confusion Analysis
For the amorphous classification task where MTL succeeds, Figure 4 shows confusion matrices. The key improvement is in true positive rate (recall): Structured MTL correctly identifies 72.2% of amorphous-forming alloys compared to 61.7% for independent models.
Figure 4: Confusion matrices for amorphous classification. MTL improves recall (fewer false negatives) at slight cost to precision.
This improvement is particularly relevant for materials screening, where missing a potential candidate (false negative) is typically more costly than investigating a false positive.
4.6 Ablation Studies: Diagnosing Negative Transfer
To understand the mechanisms behind negative transfer, we conducted three ablation experiments that directly test our hypotheses about data imbalance, gradient conflict, and feature transferability.
4.6.1 Experiment 1: Cost of Data Imbalance
We hypothesized that the 65:1 data imbalance (52,388 resistivity vs. 800 hardness samples) is the primary cause of negative transfer. To test this, we systematically downsampled the resistivity dataset to create progressively more balanced training sets: 1,000, 5,000, 10,000, 25,000, and 52,388 (full) samples.
Figure 5 shows the striking result: hardness prediction improves as the data becomes more balanced. With only 1,000 resistivity samples (roughly balanced with 800 hardness samples), hardness $R^{2}=0.830\pm 0.022$. As resistivity samples increase to 52,388, hardness $R^{2}$ degrades to $0.668\pm 0.095$—a 16.3% relative decrease.
Figure 5: Cost of data imbalance. Hardness $R^{2}$ decreases as the resistivity dataset grows, demonstrating that data imbalance directly causes negative transfer. Error bars show $\pm$1 std across 5 seeds.
Table 5 quantifies this effect across all data configurations:
Table 5: Hardness prediction performance vs. resistivity dataset size. Performance degrades monotonically with increasing imbalance.
Resistivity Samples Hardness $R^{2}$ Imbalance Ratio
1,000 0.830 $\pm$ 0.022 1.25:1
5,000 0.866 $\pm$ 0.017 6.25:1
10,000 0.845 $\pm$ 0.030 12.5:1
25,000 0.756 $\pm$ 0.045 31.25:1
52,388 0.668 $\pm$ 0.095 65.5:1
Conclusion: Data imbalance is confirmed as the primary driver of negative transfer in our setting.
4.6.2 Experiment 2: Gradient Conflict Analysis
We hypothesized that task gradients are conflicting, causing destructive interference during optimization. To test this, we computed the cosine similarity between task-specific gradients at each training step across 5 seeds and 200 epochs.
Figure 6 visualizes the mean gradient cosine similarities:
Figure 6: Gradient cosine similarity between task pairs. Values near zero indicate orthogonal gradients (no synergy); negative values indicate conflicting gradients.
The results confirm our hypothesis:
  • 1.
    Resistivity $\leftrightarrow$ Hardness: $\cos(\theta)=-0.0002\pm 0.028$ (slightly negative)
  • 2.
    Resistivity $\leftrightarrow$ Amorphous: $\cos(\theta)=0.002\pm 0.026$ (near zero)
  • 3.
    Hardness $\leftrightarrow$ Amorphous: $\cos(\theta)=0.002\pm 0.019$ (near zero)
The near-zero (or slightly negative) cosine similarities indicate that task gradients are essentially orthogonal—providing no synergy for joint optimization. The negative value for resistivity-hardness suggests slight gradient conflict, consistent with the observed negative transfer.
Conclusion: Task gradients do not synergize; they are orthogonal or slightly conflicting.
4.6.3 Experiment 3: Transfer Utility Assessment
Finally, we tested whether resistivity features can beneficially transfer to hardness prediction. We compared two approaches:
  1. 1.
    Pre-train & Transfer: Train backbone on resistivity (52,388 samples), then train hardness head with frozen backbone
  2. 2.
    Train from Scratch: Train backbone and hardness head together on hardness data only
Figure 7 shows the results:
Figure 7: Transfer learning comparison. Pre-training on resistivity provides no benefit over training from scratch for hardness prediction ($p=0.127$).
Transfer learning achieved $R^{2}=0.860\pm 0.011$, while training from scratch achieved $R^{2}=0.868\pm 0.007$. The difference is not statistically significant ($p=0.127$), and notably, transfer is slightly worse than scratch training.
Conclusion: Resistivity features do not transfer beneficially to hardness prediction, confirming fundamental task independence.
5 Discussion
5.1 Why MTL Failed for Regression
The significant degradation of regression performance under MTL (5.9% for resistivity, 16.6% for hardness) warrants careful analysis. We identify three contributing factors:
5.1.1 Severe Data Imbalance
The most striking characteristic of our dataset is the extreme imbalance: 52,388 resistivity samples versus only 800 hardness samples (65:1 ratio). Despite using inverse-proportion task weights ($w_{h}=65.0$), the optimization dynamics remain dominated by the resistivity task.
During each training batch, the expected number of resistivity samples is approximately $32\times(52388/54028)\approx 31$, while hardness samples appear in only $32\times(800/54028)\approx 0.5$ batches on average. This means most gradient updates are driven entirely by resistivity, with hardness receiving updates only occasionally. The task weights amplify the magnitude of hardness gradients when they occur, but cannot compensate for their infrequency.
Ablation Evidence: Our data quantity ablation (Section 4, Experiment 1) provides direct causal evidence for this hypothesis. When we artificially balanced the data by downsampling resistivity to 1,000 samples, hardness $R^{2}$ improved from 0.668 to 0.830—a 16.3% improvement. This demonstrates that data imbalance alone accounts for the majority of the observed negative transfer.
5.1.2 Task Independence
The learned task relation graph (Figure 1) reveals near-zero inter-task weights, indicating that predicting one property provides essentially no useful information for predicting others. This contradicts the common assumption that alloy properties are “coupled” through shared physics.
We hypothesize that while these properties do share some underlying physics (e.g., electronic structure affects both resistivity and hardness), the mapping from composition to property may not share the same functional form. Resistivity is dominated by electron scattering mechanisms, hardness by dislocation dynamics, and amorphous formation by thermodynamic stability—each governed by different combinations of atomic and structural factors.
Ablation Evidence: Our transfer utility experiment (Section 4, Experiment 3) directly tests whether resistivity features transfer to hardness prediction. The result is definitive: pre-training on 52,388 resistivity samples provides no benefit over training from scratch on hardness alone ($R^{2}=0.860$ vs. $0.868$, $p=0.127$). This confirms that learned resistivity representations do not encode useful information for hardness prediction.
5.1.3 Representation Conflict
In hard parameter sharing MTL, the shared backbone must learn a representation useful for all tasks. When tasks have conflicting optimal representations, the backbone settles into a compromise that is suboptimal for each individual task [standley2020tasks].
Given the task independence revealed by our analysis, the shared representation learned for resistivity prediction may actually be detrimental for hardness prediction, and vice versa. The backbone cannot specialize for each task, resulting in degraded performance for all.
Ablation Evidence: Our gradient conflict analysis (Section 4, Experiment 2) reveals that task gradients are nearly orthogonal ($\cos(\theta)\approx 0$) or slightly negative ($\cos(\theta)=-0.0002$ for resistivity-hardness). This means gradient updates for one task provide zero (or negative) benefit to other tasks. The slightly negative cosine between resistivity and hardness gradients directly explains why the shared backbone cannot simultaneously optimize both tasks.
5.2 Why MTL Succeeded for Classification
In contrast to regression, MTL provided significant improvements for amorphous classification, particularly in recall (+17%). We attribute this to several factors:
5.2.1 Regularization Effects
Classification tasks, especially with limited data (840 samples), are prone to overfitting. MTL provides implicit regularization through the shared backbone, which is constrained to produce representations useful across tasks. This regularization appears beneficial for the classification task, even if the inter-task information transfer is minimal.
5.2.2 Different Optimization Landscapes
Classification with cross-entropy loss has different optimization properties than regression with MSE loss. The classification loss is more robust to the gradient interference caused by multi-task optimization, allowing the model to find solutions that perform well on classification despite suboptimal regression performance.
5.2.3 Class-Level Sensitivity
The key improvement is in recall (0.722 vs. 0.617)—the model’s ability to correctly identify amorphous-forming alloys. MTL appears to improve sensitivity to the minority class (amorphous), possibly because the additional training signal from regression tasks provides useful gradients that help the classifier learn more robust decision boundaries.
5.3 Implications for Materials Discovery
Our findings have important practical implications:
5.3.1 Recommendation 1: Use Independent Models for Precise Property Prediction
When the goal is accurate prediction of specific properties (e.g., for materials selection or design), independent models are preferred. MTL provides no benefit and may actively harm performance. This is particularly true when:
  • 1.
    Properties are measured on non-overlapping sample sets
  • 2.
    Severe data imbalance exists across properties
  • 3.
    High prediction accuracy is required
5.3.2 Recommendation 2: Use MTL for Screening and Candidate Selection
When the goal is screening large candidate pools to identify promising materials, MTL can be valuable. The 17% improvement in recall means significantly fewer promising candidates are missed. In high-throughput screening scenarios, this translates directly to more successful discoveries.
5.3.3 Recommendation 3: Validate MTL Assumptions Before Deployment
Our results challenge the assumption that related material properties automatically benefit from joint learning. We recommend:
  • 1.
    Always compare MTL against independent baselines
  • 2.
    Analyze learned task relations to assess true task compatibility
  • 3.
    Consider data imbalance as a primary factor in MTL effectiveness
5.4 Limitations and Future Work
5.4.1 Dataset Limitations
Our analysis is limited to one dataset (AI-Hub Korea metal alloys) with one specific set of properties. Different property combinations or different material systems may exhibit stronger inter-task relationships. Future work should validate our findings across diverse materials datasets.
5.4.2 Gradient Balancing Methods
We used simple inverse-proportion task weighting, which proved insufficient to overcome severe imbalance. More sophisticated methods like GradNorm [chen2018gradnorm] or uncertainty weighting [kendall2018multi] may yield better results. However, our finding of task independence suggests that even perfect gradient balancing may not enable positive transfer.
5.4.3 Curriculum Learning
An alternative approach is curriculum learning: first training on the large resistivity dataset to learn general alloy representations, then fine-tuning on smaller datasets. This avoids the interference issues of simultaneous multi-task training while potentially leveraging shared knowledge.
5.4.4 Soft Parameter Sharing
Our MTL architectures used hard parameter sharing (fully shared backbone). Soft parameter sharing approaches, which allow different tasks to have different parameters with regularization encouraging similarity, may be more robust to task dissimilarity. This remains an avenue for future investigation.
6 Conclusion
This study critically examined the effectiveness of multi-task learning for metal alloy property prediction, challenging the common assumption that MTL universally improves performance when properties share underlying physics.
Using a comprehensive dataset of 54,028 metal alloy samples with three target properties (electrical resistivity, Vickers hardness, and amorphous-forming ability), we compared independent single-task models, standard MTL with hard parameter sharing, and structured MTL with learned task relation graphs. Our key findings are:
  1. 1.
    MTL significantly degrades regression performance. For resistivity prediction, $R^{2}$ dropped from 0.897 to 0.844 ($p<0.01$). For hardness, the degradation was more severe: $R^{2}$ dropped from 0.832 to 0.694 ($p<0.01$).
  2. 2.
    MTL significantly improves classification performance. For amorphous-forming ability prediction, F1 score improved from 0.703 to 0.744 ($p<0.05$), with recall improving by 17% ($p<0.05$).
  3. 3.
    Learned task relations reveal property independence. The structured MTL model learned near-zero inter-task weights ($\sim$0.006), indicating that these alloy properties do not share sufficient common structure to benefit from joint learning.
  4. 4.
    Data imbalance is a critical factor. The 65:1 ratio between resistivity and hardness samples causes the dominant task to overwhelm minority tasks during optimization, contributing to negative transfer.
These findings provide important guidelines for practitioners in materials informatics:
  • 1.
    Use independent models when precise property prediction is required
  • 2.
    Use MTL for classification/screening tasks where improved recall is valuable
  • 3.
    Validate MTL assumptions empirically rather than assuming positive transfer
  • 4.
    Consider data imbalance as a primary factor in MTL effectiveness
Future work should investigate gradient balancing methods, curriculum learning approaches, and soft parameter sharing architectures to potentially unlock the benefits of MTL even when tasks are weakly related. Additionally, expanding this analysis to other material systems and property combinations will help establish when MTL is truly beneficial in materials discovery.
Our work contributes to a more nuanced understanding of multi-task learning in materials science, moving beyond the assumption that “more data is always better” to recognize that how tasks are combined matters as much as the data itself. By providing both positive and negative results with rigorous statistical analysis, we hope to help researchers make more informed decisions about when to apply MTL in their materials discovery pipelines.