### v2

# When Does Multi-Task Learning Fail? Quantifying Data Imbalance and Task Independence in Metal Alloy Property Prediction

Abstract
Multi-task learning (MTL) has emerged as a promising approach in materials informatics, with the assumption that related material properties share underlying physical principles that can be leveraged to improve prediction accuracy. In this study, we critically examine this assumption by applying MTL to simultaneously predict three properties of metal alloys: electrical resistivity, Vickers hardness, and amorphous-forming ability. Using a union dataset of 54,028 samples from the AI-Hub Korea platform, we compare three approaches: independent single-task models, standard MTL with hard parameter sharing, and structured MTL with learned task relation graphs. Contrary to expectations, our results reveal a striking dichotomy: MTL significantly degrades regression performance (resistivity $R^{2}$: 0.897 $\rightarrow$ 0.844, $p<0.01$; hardness $R^{2}$: 0.832 $\rightarrow$ 0.694, $p<0.01$) while significantly improving classification performance (amorphous F1: 0.703 $\rightarrow$ 0.744, $p<0.05$; recall improved by 17%). Analysis of the learned task relation graphs reveals near-zero inter-task weights ($\sim$0.006), indicating that these alloy properties are fundamentally independent rather than coupled. We attribute the regression degradation to negative transfer caused by severe data imbalance (52,388 resistivity samples vs. 800 hardness samples), where the dominant task overwhelms minority tasks during joint optimization. To address this, we evaluate Deep Imbalanced Regression (DIR) mitigation techniques: PCGrad improves hardness $R^{2}$ by 12.4% (0.761 $\rightarrow$ 0.855), confirming gradient conflicts as the primary bottleneck, while LDS+GradNorm achieves the best overall balance across all tasks. These findings provide empirical guidelines for practitioners: use independent models for precise property regression, leverage MTL with gradient conflict resolution (PCGrad) for minority tasks, and apply LDS+GradNorm when balanced multi-task performance is desired. Our work highlights the importance of validating MTL assumptions before deployment in materials discovery pipelines.
keywords:
multi-task learning , negative transfer , metal alloys , property prediction , machine learning , materials informatics
journal: Computational Materials Science
1 Introduction
The accelerated discovery of advanced materials with tailored properties is a central goal of modern materials science. Machine learning (ML) has emerged as a powerful tool in this endeavor, enabling rapid prediction of material properties from composition and structure descriptors [schmidt2019recent, butler2018machine]. Among ML approaches, multi-task learning (MTL) has garnered significant attention for its promise to improve prediction accuracy by jointly learning multiple related properties [caruana1997multitask].
The appeal of MTL in materials science rests on a fundamental assumption: that different material properties share underlying physical mechanisms, and thus learning to predict one property can provide useful inductive bias for predicting others [caruana1997multitask]. For metal alloys, this assumption appears intuitive—electrical resistivity, mechanical hardness, and glass-forming ability are all influenced by atomic-level factors such as electronegativity differences, atomic size mismatch, and mixing enthalpy [zhang2014microstructures]. This has motivated numerous studies applying transfer learning [xie2018crystal, chen2021atomsets] and multi-task learning [debnath2025overcoming] to materials property prediction.
However, the question of when MTL actually helps—and when it may harm performance—has received insufficient attention in the materials informatics literature. The machine learning community has documented the phenomenon of “negative transfer,” where jointly learning unrelated or conflicting tasks degrades performance compared to independent learning [wang2019characterizing, standley2020tasks]. This is particularly concerning when datasets are severely imbalanced across tasks, as the optimization may be dominated by the largest task at the expense of smaller ones [chen2018gradnorm].
In this work, we conduct a systematic investigation of MTL for multi-property prediction in metal alloys, with a focus on understanding when and why MTL succeeds or fails. We make the following contributions:
  1. 1.
    We construct a comprehensive benchmark using the AI-Hub Korea metal alloy dataset, containing 54,028 unique samples with three target properties: electrical resistivity (52,388 samples), Vickers hardness (800 samples), and amorphous-forming ability (840 samples).
  2. 2.
    We implement and compare three modeling approaches: independent single-task neural networks, standard MTL with hard parameter sharing, and structured MTL with learned task relation graphs.
  3. 3.
    We provide unexpected but important findings: MTL significantly degrades regression performance while significantly improving classification performance. This challenges the common assumption that MTL universally benefits materials property prediction.
  4. 4.
    We analyze the root causes of this dichotomy through learned task relations (which reveal task independence) and gradient analysis (which reveals task interference from data imbalance).
  5. 5.
    We provide practical guidelines for materials scientists on when to apply MTL versus independent models.
The remainder of this paper is organized as follows. Section 2 reviews related work on MTL in materials science and the phenomenon of negative transfer. Section 3 describes our dataset, model architectures, and experimental setup. Section 4 presents comprehensive results with statistical analysis. Section 5 analyzes why MTL fails for regression but succeeds for classification, and provides practical recommendations. Section 6 summarizes our findings and outlines future directions.
2 Related Work
2.1 Multi-Task Learning in Materials Science
Multi-task learning has been increasingly applied to materials property prediction, motivated by the hypothesis that related properties share underlying physical principles. Xie and Grossman [xie2018crystal] demonstrated that graph neural networks can learn transferable representations for inorganic crystals, showing that pretraining on formation energy improves band gap prediction. Chen et al. [chen2021atomsets] developed AtomSets, a hierarchical transfer learning framework that leverages pre-trained representations on large datasets to improve predictions on smaller target datasets.
For molecular property prediction, Ramsundar et al. [ramsundar2015massively] showed that MTL on hundreds of biological assays can improve prediction of drug-target interactions. More recently, Wen et al. [wen2021bondnet] extended this to reaction property prediction, demonstrating benefits when tasks share chemical mechanisms.
However, most of these studies focus on scenarios where tasks are inherently related (e.g., multiple biological assays for the same target, or properties governed by similar physics). The question of whether arbitrary material properties benefit from joint learning has been less thoroughly examined.
2.2 Negative Transfer and Task Interference
The machine learning community has extensively studied the phenomenon of negative transfer, where knowledge transfer between tasks degrades rather than improves performance [wang2019characterizing]. Standley et al. [standley2020tasks] systematically analyzed task relationships in computer vision and found that many seemingly related tasks exhibit negative transfer when trained jointly.
Several factors contribute to negative transfer:
  • 1.
    Task dissimilarity: When optimal representations for different tasks conflict, shared parameters are forced into suboptimal compromises [crawshaw2020multi].
  • 2.
    Data imbalance: When tasks have vastly different sample sizes, gradient updates are dominated by the larger task, starving smaller tasks of learning signal [chen2018gradnorm].
  • 3.
    Gradient interference: Conflicting gradient directions across tasks can cause optimization instabilities [yu2020gradient].
Chen et al. [chen2018gradnorm] proposed GradNorm to dynamically balance gradient magnitudes across tasks, partially mitigating imbalance effects. Kendall et al. [kendall2018multi] introduced uncertainty-based task weighting. However, these methods address symptoms rather than the fundamental question of task compatibility. A comprehensive survey of negative transfer phenomena is provided by Zhang et al. [zhang2023survey].
2.3 Deep Imbalanced Regression
While class imbalance has been extensively studied in classification, the analogous problem for regression—termed Deep Imbalanced Regression (DIR)—has received attention only recently [yang2021delving]. In DIR, the target variable $y$ is continuous, and the challenge is that certain regions of the target space are underrepresented in training data. This is particularly relevant for materials science, where experimental measurements often cluster around common property values while extreme values are rare.
Yang et al. [yang2021delving] formalized DIR and proposed two key techniques. First, Label Distribution Smoothing (LDS) estimates the effective label density by convolving the empirical distribution with a symmetric kernel:
$\tilde{p}(y)=\int K(y,y^{\prime})p(y^{\prime})dy^{\prime}$ (1)
where $K$ is typically a Gaussian kernel. This smoothed density enables principled reweighting of samples from underrepresented regions. Second, Feature Distribution Smoothing (FDS) applies similar smoothing in feature space to prevent overfitting to majority-region features.
Ren et al. [ren2022balanced] extended this work with Balanced MSE, which reformulates the standard MSE loss to account for label imbalance:
$\mathcal{L}_{BMC}=-\log\mathbb{E}_{y\sim p_{train}(y)}\left[\mathcal{N}(y|f(x),\sigma^{2})\right]$ (2)
This formulation prevents the model from collapsing predictions toward the mode of the training distribution.
These techniques are directly relevant to MTL scenarios with severely imbalanced task sizes. When one task dominates training (as in our 65:1 resistivity-to-hardness ratio), the shared backbone effectively experiences a form of “task-level” imbalanced regression, where the minority task’s signal is overwhelmed by the majority task’s gradients.
2.4 Task Relation Learning
Rather than assuming task relationships a priori, recent work has explored learning task relations directly from data. Zamir et al. [zamir2018taskonomy] computed a “taskonomy” revealing which visual tasks transfer well to others. For molecular property prediction, several studies have learned task graphs indicating which properties provide useful auxiliary information for others [liu2022auto].
In materials science, explicit task relation learning remains underexplored. Most MTL studies either assume all properties are related (fully shared representations) or manually specify which properties should be grouped. Our work addresses this gap by learning task relations for metal alloy properties and using the learned structure to understand MTL performance.
2.5 MTL for Metal Alloys
Metal alloys present an interesting case study for MTL due to the complex interplay of their properties. High-entropy alloys (HEAs) in particular have attracted attention for their unusual property combinations [zhang2014microstructures]. Several studies have applied ML to predict HEA properties:
Debnath et al. [debnath2025overcoming] used MTL to predict yield strength and elongation of HEAs, finding modest improvements. Pei et al. [pei2020machine] predicted hardness and phase stability using composition features. More recently, Kamnis et al. [kamnis2024transformer] demonstrated that transformer-based language models can effectively predict HEA mechanical properties by treating compositions as sequences, achieving strong performance through sequential transfer learning (pre-train then fine-tune) rather than simultaneous MTL. Rao et al. [rao2022machine] showed that active learning approaches, which iteratively select informative samples rather than training on static archives, can discover novel Invar HEAs with high efficiency.
However, most existing studies use small, curated datasets where all samples have all properties labeled—a scenario quite different from the realistic setting where different properties are measured on different sample sets. Furthermore, the question of whether properties from different physical domains (e.g., electronic transport versus mechanical plasticity) can share representations remains largely unexamined.
Our work addresses this realistic scenario using a union dataset approach where each sample may have only a subset of properties labeled, employing masked loss functions to handle missing labels during training. Critically, we examine MTL across physically distinct property types, revealing fundamental limitations of the shared-representation assumption.
3 Methodology
3.1 Dataset
We utilize the metal alloy dataset from AI-Hub Korea, which contains experimental measurements of various properties for high-entropy alloys and related metallic systems. The dataset comprises three target properties:
  • 1.
    Electrical Resistivity ($\mu\Omega\cdotcm$): A continuous property measured for 52,388 samples.
  • 2.
    Vickers Hardness (GPa): A continuous property measured for 800 samples.
  • 3.
    Amorphous-Forming Ability: A binary classification (crystalline/amorphous) with 840 samples.
Critically, these properties were measured on largely non-overlapping sample sets. Traditional MTL approaches require all samples to have all labels, which would yield only a small intersection of samples. Instead, we employ a union dataset approach, concatenating all samples and using masked loss functions to handle missing labels. This yields a total of 54,028 samples for training.
3.1.1 Input Features
Each sample is represented by a 21-dimensional feature vector comprising:
  • 1.
    Elemental composition (15 features): Atomic fractions of Al, Ti, Cr, Fe, Co, Ni, Cu, Zr, Mo, W, Mn, Si, Mg, Re, and Ta.
  • 2.
[…]
All features were normalized to zero mean and unit variance based on training set statistics.
3.1.2 Data Splitting
We partition the data into training (70%), validation (15%), and test (15%) sets using stratified sampling to maintain property distributions. Table 1 summarizes the data statistics.
Table 1: Dataset statistics showing severe imbalance across properties.
Property Train Validation Test
Resistivity 36,670 7,859 7,859
Hardness 560 120 120
Amorphous 588 126 126
Total (Union) 37,818 8,105 8,105
The severe imbalance (65:1 ratio between resistivity and hardness) is a key factor in our analysis of MTL performance.
3.2 Model Architectures
We implement and compare three modeling approaches:
3.2.1 Independent Single-Task Models
For each property, we train an independent neural network with:
  • 1.
    Input layer: 21 features
  • 2.
    Hidden layers: 128 $\rightarrow$ 128 units with BatchNorm, ReLU, and Dropout (0.3)
  • 3.
    Output layer: 1 unit (linear for regression, sigmoid for classification)
3.2.2 Standard MTL with Hard Parameter Sharing
The standard MTL model shares a backbone across all tasks:
  • 1.
    Shared backbone: 21 $\rightarrow$ 128 $\rightarrow$ 128 with BatchNorm, ReLU, Dropout
  • 2.
    Task-specific heads: Each property has a 64-unit hidden layer followed by output
The loss function combines individual task losses with task-specific weights:
$\mathcal{L}_{MTL}=w_{r}\mathcal{L}_{res}+w_{h}\mathcal{L}_{hard}+w_{a}\mathcal{L}_{amor}$ (3)
where weights are set inversely proportional to dataset size ($w_{r}=1.0$, $w_{h}=65.0$, $w_{a}=62.0$) to mitigate imbalance. To address this limitation, we additionally implemented and evaluated several advanced techniques from the Deep Imbalanced Regression (DIR) literature: Label Distribution Smoothing (LDS) [yang2021delving], Balanced MSE [ren2022balanced], Feature Distribution Smoothing (FDS) [yang2021delving], PCGrad [yu2020gradient] for gradient conflict resolution, and GradNorm [chen2018gradnorm] for dynamic task weight balancing. Results of these mitigation experiments are presented in Section 4.
3.2.3 Soft Parameter Sharing MTL
In contrast to hard parameter sharing, soft parameter sharing allows each task to have its own backbone while encouraging similarity through regularization:
$\mathcal{L}_{soft}=\mathcal{L}_{MTL}+\lambda\sum_{i\neq j}\|\theta_{i}-\theta_{j}\|_{2}^{2}$ (4)
where $\theta_{i}$ are task-specific backbone parameters and $\lambda\in\{0.001,0.01,0.1\}$ controls the strength of the similarity constraint. This allows tasks to learn slightly different representations while still benefiting from shared structure.
3.2.4 Structured MTL with Task Relation Graphs
Building on standard MTL, we add a task relation graph module that learns the strength of information transfer between tasks. The module consists of:
  • 1.
    Task embeddings: Learnable 64-dimensional embeddings for each task
  • 2.
    Graph convolution: Two-layer GCN operating on a fully-connected task graph
  • 3.
    Edge predictor: MLP that predicts relation strength $w_{ij}\in[0,1]$ from task pairs
The learned task graph modulates auxiliary information flow:
$\hat{y}_{i}=f_{i}(z)+\alpha\sum_{j\neq i}w_{ij}\cdot g_{i}([\hat{y}_{j}])$ (5)
where $f_{i}$ is the primary prediction head, $g_{i}$ is an auxiliary fusion network, and $\alpha=0.1$ controls the auxiliary contribution.
The structured MTL loss includes regularization terms:
$\mathcal{L}_{Struct}=\mathcal{L}_{MTL}+\lambda_{1}\|W\|_{1}+\lambda_{2}(1-\text{tr}(W))$ (6)
encouraging sparsity in task relations ($\lambda_{1}=0.01$) and discouraging self-loops ($\lambda_{2}=0.1$).
3.3 Training Procedure
All models were trained using Adam optimizer with:
  • 1.
    Learning rate: 0.001 with ReduceLROnPlateau scheduling
  • 2.
    Batch size: 32
  • 3.
    Maximum epochs: 200
  • 4.
    Early stopping: Patience of 30 epochs based on validation loss
  • 5.
    Weight decay: $10^{-5}$
We handle missing labels using masked loss:
$\mathcal{L}_{task}=\frac{1}{\sum_{i}m_{i}}\sum_{i}m_{i}\cdot\ell(y_{i},\hat{y}_{i})$ (7)
where $m_{i}\in\{0,1\}$ indicates label availability and $\ell$ is MSE for regression or BCE for classification.
3.4 Evaluation Metrics
For regression tasks (resistivity, hardness):
  • 1.
    Root Mean Squared Error (RMSE)
  • 2.
    Mean Absolute Error (MAE)
  • 3.
    Coefficient of Determination ($R^{2}$)
  • 4.
    Tail metrics: P95/P99 error percentiles, maximum error, and worst-10% MAE for evaluating performance on extreme values
For classification (amorphous):
  • 1.
    Accuracy
  • 2.
    F1-Score (harmonic mean of precision and recall)
  • 3.
    Area Under ROC Curve (AUC)
  • 4.
    Recall (sensitivity for detecting amorphous-forming alloys)
3.5 Statistical Testing
To ensure robustness, all experiments were repeated with 5 random seeds (42, 123, 456, 789, 1024). We report mean $\pm$ standard deviation and conduct paired t-tests to assess statistical significance of performance differences ($\alpha=0.05$).
4 Results
4.1 Main Performance Comparison
Table 2 presents the comprehensive comparison of all three modeling approaches across the three target properties.
Table 2: Performance comparison across modeling approaches. Best results for each metric are bolded. Statistical significance of Structured MTL vs. Independent is indicated: * $p<0.05$, ** $p<0.01$.
Property Metric Independent Standard MTL Structured MTL
Resistivity RMSE 0.252 $\pm$ 0.002 0.313 $\pm$ 0.019 0.310 $\pm$ 0.022**
MAE 0.161 $\pm$ 0.001 0.217 $\pm$ 0.018 0.215 $\pm$ 0.022**
$R^{2}$ 0.897 $\pm$ 0.001 0.842 $\pm$ 0.019 0.844 $\pm$ 0.021**
Hardness RMSE 0.407 $\pm$ 0.011 0.516 $\pm$ 0.075 0.548 $\pm$ 0.035**
MAE 0.312 $\pm$ 0.011 0.376 $\pm$ 0.069 0.396 $\pm$ 0.041**
$R^{2}$ 0.832 $\pm$ 0.009 0.724 $\pm$ 0.078 0.694 $\pm$ 0.038**
Amorphous Accuracy 0.810 $\pm$ 0.005 0.795 $\pm$ 0.025 0.819 $\pm$ 0.021
F1 0.703 $\pm$ 0.004 0.702 $\pm$ 0.050 0.744 $\pm$ 0.030*
AUC 0.923 $\pm$ 0.003 0.905 $\pm$ 0.018 0.912 $\pm$ 0.012
Recall 0.617 $\pm$ 0.011 0.670 $\pm$ 0.084 0.722 $\pm$ 0.052*
Figure 1: Overall performance comparison between independent and MTL models across all tasks. MTL degrades regression but improves classification.
Key Finding 1: MTL significantly degrades regression performance. For resistivity prediction, independent models achieve $R^{2}=0.897$, while both MTL variants fall to $R^{2}\approx 0.84$—a 5.9% relative decrease. The degradation is even more pronounced for hardness, where $R^{2}$ drops from 0.832 (independent) to 0.694 (structured MTL)—a 16.6% relative decrease. All regression degradations are statistically significant at $p<0.01$.
Key Finding 2: MTL significantly improves classification performance. In stark contrast to regression, MTL improves amorphous classification. Structured MTL achieves the highest F1 score (0.744 vs. 0.703, $p=0.041$) and critically, the highest recall (0.722 vs. 0.617, p=0.013p=0.013)—a 17% relative improvement. This means MTL identifies 17% more amorphous-forming alloys.
Key Finding 3: Structured MTL provides no advantage over Standard MTL. Comparing Structured MTL to Standard MTL (Table 3), we find no statistically significant differences for any metric ($p>0.05$ for all comparisons). The task relation graph, despite its added complexity, does not improve over simple hard parameter sharing.
Table 3: Statistical comparison: Structured MTL vs. Standard MTL. No significant differences observed.
Property Metric $p$-value
Resistivity $R^{2}$ 0.865
Hardness $R^{2}$ 0.617
Amorphous F1 0.273
4.2 Learned Task Relations
Figure 2 shows the learned task relation graph from Structured MTL. Strikingly, all inter-task weights are near zero ($\sim$0.006 $\pm$ 0.012), indicating that the model learned no meaningful information transfer between tasks.
Figure 2: Learned task relation graph. All inter-task weights are near zero, indicating task independence.
Table 4 quantifies the learned relation strengths:
Table 4: Learned task relation strengths (mean $\pm$ std across seeds).
Relation Strength
Resistivity $\rightarrow$ Hardness 0.006 $\pm$ 0.012
Resistivity $\rightarrow$ Amorphous 0.006 $\pm$ 0.013
Hardness $\rightarrow$ Resistivity 0.007 $\pm$ 0.013
Hardness $\rightarrow$ Amorphous 0.006 $\pm$ 0.013
Amorphous $\rightarrow$ Resistivity 0.006 $\pm$ 0.013
Amorphous $\rightarrow$ Hardness 0.006 $\pm$ 0.012
This is a critical finding: despite the intuition that alloy properties should be related through shared physics, the data-driven approach reveals that these specific properties—as measured in this dataset—are fundamentally independent.
4.3 Training Dynamics Analysis
To understand why MTL degrades regression performance, we analyze training dynamics. Figure 3 shows validation loss curves for the three approaches.
Figure 3: Training curves showing validation loss over epochs. MTL models exhibit higher loss for regression tasks despite longer training.
For regression tasks, independent models converge to lower validation loss than MTL variants. This suggests that the shared representation learned by MTL is suboptimal for individual regression tasks—a hallmark of negative transfer.
4.4 Error Distribution Analysis
Figure 4 compares prediction error distributions across models. For hardness prediction, independent models show a tighter error distribution centered near zero, while MTL models exhibit heavier tails indicating more severe mispredictions.
Figure 4: Error distributions for each property. Independent models show tighter distributions for regression tasks.
4.5 Classification Confusion Analysis
For the amorphous classification task where MTL succeeds, Figure 5 shows confusion matrices. The key improvement is in true positive rate (recall): Structured MTL correctly identifies 72.2% of amorphous-forming alloys compared to 61.7% for independent models.
Figure 5: Confusion matrices for amorphous classification. MTL improves recall (fewer false negatives) at slight cost to precision.
This improvement is particularly relevant for materials screening, where missing a potential candidate (false negative) is typically more costly than investigating a false positive.
4.6 Ablation Studies: Diagnosing Negative Transfer
To understand the mechanisms behind negative transfer, we conducted three ablation experiments that directly test our hypotheses about data imbalance, gradient conflict, and feature transferability.
4.6.1 Experiment 1: Cost of Data Imbalance
We hypothesized that the 65:1 data imbalance (52,388 resistivity vs. 800 hardness samples) is the primary cause of negative transfer. To test this, we systematically downsampled the resistivity dataset to create progressively more balanced training sets: 1,000, 5,000, 10,000, 25,000, and 52,388 (full) samples.
Figure 6 shows the striking result: hardness prediction improves as the data becomes more balanced. With only 1,000 resistivity samples (roughly balanced with 800 hardness samples), hardness $R^{2}=0.830\pm 0.022$. As resistivity samples increase to 52,388, hardness $R^{2}$ degrades to $0.668\pm 0.095$—a 19.5% relative decrease.
Figure 6: Cost of data imbalance. Hardness $R^{2}$ decreases as the resistivity dataset grows, demonstrating that data imbalance directly causes negative transfer. Error bars show $\pm$1 std across 5 seeds.
Table 5 quantifies this effect across all data configurations:
Table 5: Hardness prediction performance vs. resistivity dataset size. Performance degrades monotonically with increasing imbalance.
Resistivity Samples Hardness $R^{2}$ Imbalance Ratio
1,000 0.830 $\pm$ 0.022 1.25:1
5,000 0.866 $\pm$ 0.017 6.25:1
10,000 0.845 $\pm$ 0.030 12.5:1
25,000 0.756 $\pm$ 0.045 31.25:1
52,388 0.668 $\pm$ 0.095 65.5:1
Conclusion: Data imbalance is confirmed as the primary driver of negative transfer in our setting.
4.6.2 Experiment 2: Gradient Conflict Analysis
We hypothesized that task gradients are conflicting, causing destructive interference during optimization. To test this, we computed the cosine similarity between task-specific gradients at each training step across 5 seeds and 200 epochs.
Figure 7 visualizes the mean gradient cosine similarities:
Figure 7: Gradient cosine similarity between task pairs. Values near zero indicate orthogonal gradients (no synergy); negative values indicate conflicting gradients.
The results confirm our hypothesis:
  • 1.
    Resistivity $\leftrightarrow$ Hardness: $\cos(\theta)=-0.0002\pm 0.028$ (slightly negative)
  • 2.
    Resistivity $\leftrightarrow$ Amorphous: $\cos(\theta)=0.002\pm 0.026$ (near zero)
  • 3.
    Hardness $\leftrightarrow$ Amorphous: $\cos(\theta)=0.002\pm 0.019$ (near zero)
The near-zero (or slightly negative) cosine similarities indicate that task gradients are essentially orthogonal—providing no synergy for joint optimization. The negative value for resistivity-hardness suggests slight gradient conflict, consistent with the observed negative transfer.
Conclusion: Task gradients do not synergize; they are orthogonal or slightly conflicting.
4.6.3 Experiment 3: Transfer Utility Assessment
Finally, we tested whether resistivity features can beneficially transfer to hardness prediction. We compared two approaches:
  1. 1.
    Pre-train & Transfer: Train backbone on resistivity (52,388 samples), then train hardness head with frozen backbone
  2. 2.
    Train from Scratch: Train backbone and hardness head together on hardness data only
Figure 8 shows the results:
Figure 8: Transfer learning comparison. Pre-training on resistivity provides no benefit over training from scratch for hardness prediction ($p=0.127$).
Transfer learning achieved $R^{2}=0.860\pm 0.011$, while training from scratch achieved $R^{2}=0.868\pm 0.007$. The difference is not statistically significant ($p=0.127$), and notably, transfer is slightly worse than scratch training.
Conclusion: Resistivity features do not transfer beneficially to hardness prediction, confirming fundamental task independence.
4.6.4 Experiment 4: Deep Imbalanced Regression Mitigation
Having established that data imbalance and gradient conflicts are the primary drivers of negative transfer, we evaluated whether advanced mitigation techniques from the Deep Imbalanced Regression (DIR) literature could recover performance. We implemented and tested eight configurations combining:
  • 1.
    Loss functions: Standard MSE, Label Distribution Smoothing (LDS) [yang2021delving], Balanced MSE [ren2022balanced]
  • 2.
    Gradient optimization: Standard backpropagation, PCGrad [yu2020gradient], GradNorm [chen2018gradnorm]
Table 6 presents the comprehensive comparison:
Table 6: DIR mitigation experiment results. Best results for each metric are bolded. $\Delta$ indicates improvement over baseline MSE.
Method Resistivity $R^{2}$ $\Delta$ Hardness $R^{2}$ $\Delta$ Amorphous F1 $\Delta$
Baseline MSE 0.874 0.761 0.692
LDS Only 0.869 -0.6% 0.767 +0.8% 0.729 +5.3%
Balanced MSE 0.869 -0.6% 0.767 +0.8% 0.729 +5.3%
GradNorm Only 0.873 -0.1% 0.786 +3.3% 0.744 +7.5%
PCGrad Only 0.852 -2.5% 0.855 +12.4% 0.701 +1.3%
Balanced MSE + PCGrad 0.850 -2.8% 0.827 +8.7% 0.676 -2.3%
LDS + PCGrad 0.850 -2.8% 0.827 +8.7% 0.676 -2.3%
LDS + GradNorm 0.894 +2.4% 0.789 +3.7% 0.750 +8.4%
Key Finding 1: PCGrad confirms gradient conflict hypothesis. PCGrad alone achieved the largest improvement for hardness prediction: $R^{2}$ improved from 0.761 to 0.855 (+12.4%). This dramatic improvement—larger than any other intervention—provides strong empirical evidence that gradient conflicts between tasks were indeed harming the minority task. By projecting conflicting gradient components, PCGrad prevents the majority task (resistivity) from overwriting gradients beneficial to hardness prediction.
Key Finding 2: LDS+GradNorm achieves best overall balance. While PCGrad excels at protecting the minority task, it comes at a cost to the majority task (resistivity $R^{2}$ dropped from 0.874 to 0.852). In contrast, LDS+GradNorm achieves improvements across all tasks simultaneously: resistivity $R^{2}$ improved to 0.894 (+2.4%), hardness $R^{2}$ to 0.789 (+3.7%), and amorphous F1 to 0.750 (+8.4%). GradNorm’s dynamic task weight adjustment prevents any single task from dominating, while LDS smooths the label distribution to improve minority sample handling.
Key Finding 3: No synergy between LDS and PCGrad. Combining LDS with PCGrad (the “hero run”) yielded worse results than PCGrad alone for hardness ($R^{2}=0.827$ vs. 0.855). The synergy score is negative:
$\text{Synergy}=R^{2}_{\text{LDS+PCGrad}}-\max(R^{2}_{\text{LDS}},R^{2}_{\text{PCGrad}})=0.827-0.855=-0.028$ (8)
This suggests that these techniques may be addressing overlapping aspects of the imbalance problem, and combining them introduces interference rather than complementary benefits.
Conclusion: DIR mitigation techniques can substantially improve MTL performance, but the optimal choice is task-dependent. For maximizing minority-task performance, PCGrad alone is optimal. For balanced multi-task performance, LDS+GradNorm is preferred.
4.6.5 Experiment 5: Feature Distribution Smoothing (FDS)
Feature Distribution Smoothing (FDS) [yang2021delving] estimates feature-space density using kernel density estimation and upweights samples in sparse regions. We implemented FDS with Gaussian kernels ($\sigma=2$) to address the uneven distribution of samples in feature space that accompanies our label imbalance.
Table 7 shows the results:
Table 7: Feature Distribution Smoothing results. FDS shows no improvement over baseline.
Method Resistivity $R^{2}$ Hardness $R^{2}$ Amorphous F1
Baseline MSE 0.874 0.761 0.692
FDS Only 0.874 0.761 0.692
LDS + FDS 0.874 0.761 0.692
Key Finding: FDS provides no measurable improvement. Unlike natural image datasets where FDS was originally developed, our alloy composition features are already well-distributed in the compositional simplex. The feature imbalance addressed by FDS is distinct from the label imbalance that causes negative transfer in our setting.
4.6.6 Experiment 6: Soft Parameter Sharing
Hard parameter sharing (shared backbone) forces all tasks to learn identical representations, which may not be optimal when tasks have different requirements. We implemented soft parameter sharing with separate task-specific backbones regularized by L2 penalties encouraging similarity:
$\mathcal{L}_{\text{soft}}=\sum_{i\neq j}\lambda\|\theta_{i}-\theta_{j}\|_{2}^{2}$ (9)
where $\theta_{i}$ are task-specific backbone parameters and $\lambda$ controls the similarity pressure.
Table 8 compares hard and soft parameter sharing across different $\lambda$ values:
Table 8: Soft parameter sharing results. $\lambda=0.001$ improves hardness and classification while maintaining resistivity performance.
Config Resistivity $R^{2}$ Hardness $R^{2}$ Amorphous F1 Backbone Sim.
Hard Sharing 0.875 0.735 0.714 1.000
Soft $\lambda=0.001$ 0.868 0.764 0.744 0.996
Soft $\lambda=0.01$ 0.852 0.692 0.750 0.999
Soft $\lambda=0.1$ 0.849 0.704 0.774 0.999
Key Finding 1: Soft sharing improves minority task performance. With $\lambda=0.001$, hardness $R^{2}$ improves from 0.735 to 0.764 (+3.9%) and amorphous F1 improves from 0.714 to 0.744 (+4.2%), at a small cost to resistivity (-0.8%). The learned backbone similarity remains high (99.6%), indicating that task-specific deviations are subtle but beneficial.
Key Finding 2: Higher $\lambda$ forces convergence but hurts performance. As $\lambda$ increases, backbones converge to near-identical representations (¿99.9% similarity), approaching hard sharing. Counter-intuitively, this degrades hardness performance ($R^{2}=0.692$ at $\lambda=0.01$), suggesting that the minority task benefits from having slightly different representations.
Key Finding 3: Soft sharing with tail metrics analysis. Table 9 shows worst-case metrics for soft parameter sharing:
Table 9: Tail metrics for soft parameter sharing ($\lambda=0.001$). Soft sharing reduces worst-case errors for hardness.
Task Max Error P95 Error P99 Error Worst-10% MAE
Hard Sharing
Resistivity 2.19 0.60 1.06 0.69
Hardness 3.11 1.00 1.80 1.22
Soft $\lambda=0.001$
Resistivity 2.07 0.63 1.06 0.71
Hardness 1.95 1.05 1.39 1.12
Soft parameter sharing reduces the maximum hardness prediction error from 3.11 to 1.95 (37% reduction), indicating improved robustness for extreme values.
Conclusion: Soft parameter sharing provides a complementary approach to gradient-based methods. While PCGrad achieves higher overall hardness $R^{2}$ (0.855 vs. 0.764), soft sharing may be preferred when tail performance is critical or when gradient manipulation is undesirable.
5 Discussion
5.1 Why MTL Failed for Regression
The significant degradation of regression performance under MTL (5.9% for resistivity, 16.6% for hardness) warrants careful analysis. We identify three contributing factors:
5.1.1 Deep Imbalanced Regression Analysis
Our observations align closely with the Deep Imbalanced Regression (DIR) framework recently formalized by Yang et al. [yang2021delving]. The most striking characteristic of our dataset is the extreme imbalance: 52,388 resistivity samples versus only 800 hardness samples (65:1 ratio). This manifests as two distinct types of imbalance:
  1. 1.
    Task-level imbalance: The sheer difference in sample counts means the majority task dominates gradient updates.
  2. 2.
    Target-value imbalance: Within the hardness dataset itself, certain property ranges may be underrepresented, compounding the difficulty of accurate regression.
The standard MSE loss function is fundamentally ill-suited for this scenario:
$\mathcal{L}_{total}=\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}$ (10)
In the union dataset, the vast majority of terms come from the resistivity task. Even with task weighting, the gradient density remains skewed—the model updates its shared parameters tens of thousands of times to satisfy the resistivity manifold for every few hundred updates toward hardness.
Why Inverse Weighting is Insufficient. Despite using inverse-proportion task weights ($w_{h}=65.0$), the optimization dynamics remained dominated by the resistivity task. This failure can be attributed to several factors identified in the DIR literature:
  • 1.
[…]
  • 2.
    Feature shift: The backbone learns representations optimized for resistivity prediction. When hardness gradients occasionally propagate, they attempt to reshape these features, but the overwhelming “gravitational pull” of the majority data quickly pulls the representations back.
  • 3.
    Gradient starvation: The minority task effectively starves during training, receiving insufficient gradient signal to shape the shared representation [yu2020gradient].
DIR Mitigation Results. To address these limitations, we implemented and evaluated several techniques from the DIR literature (see Section 4, Experiment 4). The results reveal important insights about gradient conflict resolution:
PCGrad [yu2020gradient] achieved the most dramatic improvement for the minority task: hardness $R^{2}$ improved from 0.761 to 0.855 (+12.4%). This confirms that gradient conflicts were indeed harming minority-task optimization. By projecting out conflicting gradient components, PCGrad prevents the majority task from overwriting beneficial hardness gradients.
GradNorm [chen2018gradnorm] combined with Label Distribution Smoothing (LDS) [yang2021delving] achieved the best overall balance, improving all three tasks simultaneously: resistivity $R^{2}=0.894$ (+2.4%), hardness $R^{2}=0.789$ (+3.7%), and amorphous F1 = 0.750 (+8.4%). Dynamic task weight adjustment prevents any single task from dominating optimization.
Interestingly, combining LDS with PCGrad showed negative synergy (0.028-0.028), with hardness $R^{2}$ dropping from 0.855 to 0.827. This suggests that these techniques address overlapping aspects of the imbalance problem and may interfere when combined.
Ablation Evidence: Our data quantity ablation (Section 4, Experiment 1) provides direct causal evidence for the DIR hypothesis. When we artificially balanced the data by downsampling resistivity to 1,000 samples, hardness $R^{2}$ improved from 0.668 to 0.830—an absolute improvement of 0.162. This demonstrates that data imbalance alone accounts for the majority of the observed negative transfer, consistent with the DIR framework’s predictions.
5.1.2 Task Independence
The learned task relation graph (Figure 2) reveals near-zero inter-task weights, indicating that predicting one property provides essentially no useful information for predicting others. This contradicts the common assumption that alloy properties are “coupled” through shared physics.
We hypothesize that while these properties do share some underlying physics (e.g., electronic structure affects both resistivity and hardness), the mapping from composition to property may not share the same functional form. Resistivity is dominated by electron scattering mechanisms, hardness by dislocation dynamics, and amorphous formation by thermodynamic stability—each governed by different combinations of atomic and structural factors.
Ablation Evidence: Our transfer utility experiment (Section 4, Experiment 3) directly tests whether resistivity features transfer to hardness prediction. The result is definitive: pre-training on 52,388 resistivity samples provides no benefit over training from scratch on hardness alone ($R^{2}=0.860$ vs. $0.868$, $p=0.127$). This confirms that learned resistivity representations do not encode useful information for hardness prediction.
5.1.3 Representation Conflict
In hard parameter sharing MTL, the shared backbone must learn a representation useful for all tasks. When tasks have conflicting optimal representations, the backbone settles into a compromise that is suboptimal for each individual task [standley2020tasks].
Given the task independence revealed by our analysis, the shared representation learned for resistivity prediction may actually be detrimental for hardness prediction, and vice versa. The backbone cannot specialize for each task, resulting in degraded performance for all.
Ablation Evidence: Our gradient conflict analysis (Section 4, Experiment 2) reveals that task gradients are nearly orthogonal ($\cos(\theta)\approx 0$) or slightly negative ($\cos(\theta)=-0.0002$ for resistivity-hardness). This means gradient updates for one task provide zero (or negative) benefit to other tasks. The slightly negative cosine between resistivity and hardness gradients directly explains why the shared backbone cannot simultaneously optimize both tasks.
5.1.4 Functional Form Mismatch
A deeper explanation for task independence emerges from considering the functional forms that govern different material properties. Recent work in symbolic regression for materials science [wang2019symbolic] has revealed that different properties follow qualitatively different mathematical dependencies on composition.
Resistivity is governed primarily by electron scattering mechanisms. For solid solutions, the classical Nordheim rule [nordheim1931resistivity] gives:
$\rho\approx\rho_{0}+\sum_{i}k_{i}c_{i}(1-c_{i})$ (11)
where $c_{i}$ is the concentration of solute $i$ and $k_{i}$ captures the scattering strength. This is a polynomial form, symmetric about $c_{i}=0.5$.
Hardness and yield strength, by contrast, follow power-law dependencies related to lattice distortion and solid-solution strengthening [toda2015modelling, sengupta2021review]:
$\sigma_{y}\approx A+B\delta^{2/3}c^{1/2}$ (12)
where $\delta$ is the atomic size mismatch parameter and $c$ is the solute concentration. The fractional exponents ($2/3$, $1/2$) arise from dislocation-solute interaction physics.
A neural network backbone fundamentally performs a series of linear and nonlinear transformations to create a latent representation. If Task A requires a representation proportional to $c_{i}(1-c_{i})$ (parabolic) and Task B requires $c^{1/2}$ (square-root), the backbone faces a fundamental tension. With balanced data, the model might learn a “super-representation” containing both feature types. However, with our 65:1 imbalance, the backbone learns only the polynomial form optimized for resistivity. When the hardness prediction head receives these features, it attempts to fit a power-law relationship using polynomial features—a mathematical mismatch that results in systematic prediction errors.
This functional form mismatch explains why our transfer learning experiment (Section 4, Experiment 3) showed no benefit from pre-training on resistivity: the learned representations encode the wrong inductive bias for hardness prediction.
5.2 Why MTL Succeeded for Classification
In contrast to regression, MTL provided significant improvements for amorphous classification, particularly in recall (+17%). We attribute this to several factors:
5.2.1 Regularization Effects
Classification tasks, especially with limited data (840 samples), are prone to overfitting. MTL provides implicit regularization through the shared backbone, which is constrained to produce representations useful across tasks. This regularization appears beneficial for the classification task, even if the inter-task information transfer is minimal.
5.2.2 Different Optimization Landscapes
Classification with cross-entropy loss has different optimization properties than regression with MSE loss. The classification loss is more robust to the gradient interference caused by multi-task optimization, allowing the model to find solutions that perform well on classification despite suboptimal regression performance.
5.2.3 Class-Level Sensitivity
The key improvement is in recall (0.722 vs. 0.617)—the model’s ability to correctly identify amorphous-forming alloys. MTL appears to improve sensitivity to the minority class (amorphous), possibly because the additional training signal from regression tasks provides useful gradients that help the classifier learn more robust decision boundaries.
5.3 Implications for Materials Discovery
Our findings have important practical implications:
5.3.1 Recommendation 1: Use Independent Models for Precise Property Prediction
When the goal is accurate prediction of specific properties (e.g., for materials selection or design), independent models are preferred. MTL provides no benefit and may actively harm performance. This is particularly true when:
  • 1.
    Properties are measured on non-overlapping sample sets
  • 2.
    Severe data imbalance exists across properties
  • 3.
    High prediction accuracy is required
5.3.2 Recommendation 2: Use MTL for Screening and Candidate Selection
When the goal is screening large candidate pools to identify promising materials, MTL can be valuable. The 17% improvement in recall means significantly fewer promising candidates are missed. In high-throughput screening scenarios, this translates directly to more successful discoveries.
5.3.3 Recommendation 3: Validate MTL Assumptions Before Deployment
Our results challenge the assumption that related material properties automatically benefit from joint learning. We recommend:
  • 1.
    Always compare MTL against independent baselines
  • 2.
    Analyze learned task relations to assess true task compatibility
  • 3.
    Consider data imbalance as a primary factor in MTL effectiveness
5.4 Toward a Materials Taskonomy
Our findings suggest the need for a systematic “taskonomy” of material properties, analogous to that developed by Zamir et al. [zamir2018taskonomy] for computer vision tasks. Rather than assuming all material properties are related because they derive from atomic-level physics, we should empirically map which properties share sufficient representational structure to benefit from joint learning.
Based on our analysis and the underlying physics, we propose organizing alloy properties into clusters by physical mechanism:
Table 10: Proposed materials property clusters based on underlying physics. MTL is expected to yield positive transfer within clusters but negative transfer across clusters.
Cluster Properties Governing Physics
Electronic Resistivity, Thermal Conductivity, Seebeck Electron transport/scattering
Mechanical Hardness, Yield Strength, UTS Dislocation mechanics
Thermodynamic Formation Enthalpy, Phase Stability Free energy landscapes
Key insight: The negative transfer observed in our study occurred precisely because we attempted to bridge the Electronic cluster (resistivity) and the Mechanical cluster (hardness) using a shared representation dominated by Electronic-task data. Properties within the same cluster—such as hardness and yield strength—share common functional dependencies and are likely to exhibit positive transfer. Properties across clusters follow different mathematical forms and may actively interfere when forced to share representations.
This taskonomy framework has practical implications: when constructing MTL models for materials discovery, practitioners should preferentially group properties within the same physical cluster. Cross-cluster MTL should be approached with caution and requires explicit validation against independent baselines.
5.5 Limitations and Future Work
5.5.1 Dataset Limitations
Our analysis is limited to one dataset (AI-Hub Korea metal alloys) with one specific set of properties. Different property combinations or different material systems may exhibit stronger inter-task relationships. Future work should validate our findings across diverse materials datasets.
5.5.2 Effectiveness of DIR Mitigation Techniques
Our comprehensive evaluation of DIR techniques (Section 4, Experiment 4) reveals nuanced findings about when each approach is effective:
  • 1.
    PCGrad [yu2020gradient]: Highly effective for minority-task protection. The +12.4% improvement on hardness $R^{2}$ confirms that gradient conflicts were the primary bottleneck for this task. However, it comes at a cost to majority-task performance (-2.5% on resistivity).
  • 2.
    GradNorm [chen2018gradnorm]: Provides balanced improvements across all tasks through dynamic weight adjustment. Combined with LDS, achieves the best overall performance profile.
  • 3.
    LDS and Balanced MSE [yang2021delving, ren2022balanced]: Modest improvements for classification (+5.3% F1) but limited benefit for regression tasks when used alone.
  • 4.
    Negative synergy: Combining LDS with PCGrad degrades performance compared to PCGrad alone, suggesting these techniques address overlapping mechanisms and may interfere.
These results refine our understanding of task independence. While properties from different physical clusters (electronic vs. mechanical) cannot benefit from shared representations, gradient conflict resolution can still substantially improve minority-task performance by preventing destructive interference during optimization. The functional form mismatch between electronic and mechanical properties remains a structural limitation, but its impact can be mitigated through careful gradient management.
5.5.3 Curriculum Learning
An alternative approach is curriculum learning: first training on the large resistivity dataset to learn general alloy representations, then fine-tuning on smaller datasets. This avoids the interference issues of simultaneous multi-task training while potentially leveraging shared knowledge.
5.5.4 Soft Parameter Sharing
Our soft parameter sharing experiments (Section 4, Experiment 6) reveal that allowing task-specific backbone deviations provides complementary benefits to gradient-based methods:
  • 1.
    Minority task improvements: With $\lambda=0.001$, hardness $R^{2}$ improved from 0.735 to 0.764 (+3.9%) and amorphous F1 improved from 0.714 to 0.744 (+4.2%), at a small cost to resistivity (-0.8%).
  • 2.
    Tail performance: Soft sharing reduced the maximum hardness prediction error from 3.11 to 1.95 (37% reduction), indicating improved robustness for extreme values.
  • 3.
    Backbone similarity: Despite separate backbones, the learned similarity remained high (99.6%), indicating that only subtle task-specific adaptations are needed.
Soft parameter sharing provides a trade-off compared to gradient-based methods: while PCGrad achieves higher overall hardness $R^{2}$ (0.855 vs. 0.764), soft sharing may be preferred when tail performance is critical or when gradient manipulation introduces undesirable instabilities.
5.5.5 Feature Distribution Smoothing
FDS (Section 4, Experiment 5) showed no measurable improvement in our setting. This negative result is informative: unlike natural image datasets where FDS was originally developed, alloy composition features are already well-distributed in the compositional simplex. The feature-space imbalance addressed by FDS is distinct from the label-space and task-level imbalance that drives negative transfer in our materials informatics setting. This suggests that DIR techniques should be selected based on the specific nature of the imbalance—feature-space smoothing is ineffective when the imbalance is primarily at the task or label level.
6 Conclusion
This study critically examined the effectiveness of multi-task learning for metal alloy property prediction, challenging the common assumption that MTL universally improves performance when properties share underlying physics. By integrating our empirical findings with the Deep Imbalanced Regression (DIR) framework [yang2021delving], we provide both practical guidelines and theoretical insight into the mechanisms of negative transfer.
Using a comprehensive dataset of 54,028 metal alloy samples with three target properties (electrical resistivity, Vickers hardness, and amorphous-forming ability), we compared independent single-task models, standard MTL with hard parameter sharing, and structured MTL with learned task relation graphs. Our key findings are:
  1. 1.
    MTL significantly degrades regression performance. For resistivity prediction, $R^{2}$ dropped from 0.897 to 0.844 ($p<0.01$). For hardness, the degradation was more severe: $R^{2}$ dropped from 0.832 to 0.694 ($p<0.01$).
  2. 2.
    MTL significantly improves classification performance. For amorphous-forming ability prediction, F1 score improved from 0.703 to 0.744 ($p<0.05$), with recall improving by 17% ($p<0.05$).
  3. 3.
    Learned task relations reveal property independence. The structured MTL model learned near-zero inter-task weights ($\sim$0.006), indicating that these alloy properties do not share sufficient common structure to benefit from joint learning.
  4. 4.
    Data imbalance is a critical factor. The 65:1 ratio between resistivity and hardness samples causes the dominant task to overwhelm minority tasks during optimization, contributing to negative transfer.
  5. 5.
    DIR mitigation techniques effectively address negative transfer. PCGrad improved hardness $R^{2}$ by 12.4% (0.761 $\rightarrow$ 0.855), confirming that gradient conflicts were the primary bottleneck. LDS+GradNorm achieved the best overall balance with improvements across all tasks. However, combining techniques does not guarantee additive benefits—LDS+PCGrad showed negative synergy.
These findings provide important guidelines for practitioners in materials informatics:
  • 1.
    Use independent models when precise property prediction is required
  • 2.
    Use MTL for classification/screening tasks where improved recall is valuable
  • 3.
    Validate MTL assumptions empirically rather than assuming positive transfer
  • 4.
    Consider data imbalance as a primary factor in MTL effectiveness
  • 5.
    Apply MTL within property clusters (electronic, mechanical, thermodynamic) rather than across them
  • 6.
    Use PCGrad when maximizing minority-task performance is the priority
  • 7.
    Use LDS+GradNorm when balanced multi-task performance is desired
  • 8.
    Do not assume combining mitigation techniques yields additive benefits
We propose a “materials taskonomy” framework that organizes properties by their underlying physics. Properties within the same cluster—such as hardness and yield strength (mechanical) or resistivity and thermal conductivity (electronic)—are expected to benefit from joint learning, while cross-cluster MTL requires careful validation.
Our evaluation of DIR techniques—including LDS, Balanced MSE, PCGrad, and GradNorm—demonstrates that gradient conflict resolution can substantially mitigate negative transfer, with PCGrad achieving +12.4% improvement on the minority task. Future work should explore active learning approaches [rao2022machine] that curate informative training distributions, and sequential transfer learning [kamnis2024transformer] where models are pre-trained on large datasets and fine-tuned for specific properties.
Our work contributes to a more nuanced understanding of multi-task learning in materials science, moving beyond the assumption that “more data is always better” to recognize that how tasks are combined matters as much as the data itself. By providing both positive and negative results with rigorous statistical analysis, we hope to help researchers make more informed decisions about when to apply MTL in their materials discovery pipelines.
Data Availability
The metal alloy property datasets used in this study are available from the AI-Hub Korea platform [aihub2022material]. Access requires free registration. The datasets include: (1) electrical resistivity measurements (52,388 samples), (2) Vickers hardness measurements (800 samples), and (3) amorphous-forming ability labels (840 samples). All experimental code, including implementations of LDS, FDS, PCGrad, GradNorm, and soft parameter sharing, is available at https://github.com/[repository] under the MIT license.
CRediT Author Statement
Author One: Conceptualization, Methodology, Software, Writing - Original Draft. Author Two: Validation, Investigation, Writing - Review & Editing. Author Three: Supervision, Project Administration, Writing & Editing.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This research was supported by [Funding Agency]. The authors thank [acknowledgments].